Calculate the surface area for a wall that measures 10 feet tall and 48 feet long. The wall has 1 window 3 feet by 3 feet, 2 windows 4 feet by 4 feet, 1 window 2 feet by 3 feet and 3 doors 3 feet by 7 feet. 


1
Expert's answer
2021-11-16T19:30:19-0500

The total area of the wall is


"10(48)=480(ft^2)"

The area of 1 window 3 feet by 3 feet is


"1(3)(3)=9(ft^2)"

The area of 2 windows 4 feet by 4 feet is


"2(4)(4)=32(ft^2)"

The area of 1 window 2 feet by 3 feet is


"1(2)(3)=6(ft^2)"

The area of 3 doors 3 feet by 7 feet is


"3(3)(7)=63(ft^2)"

Then the surface area for a wall is


"S=480-9-32-6-63=370(ft^2)"

The surface area for a wall is 370 square feet.